# Unit Test Generation for Vulnerability Exploitation in Java Third-Party Libraries

### v1

@@ Abstract
Open-source third-party libraries are widely used in software development. These libraries offer substantial advantages in terms of time and resource savings. However, a significant concern arises due to the publicly disclosed vulnerabilities within these libraries. Existing automated vulnerability detection tools often suffer from false positives and fail to accurately assess the propagation of inputs capable of triggering vulnerabilities from client projects to vulnerable code in libraries. In this paper, we propose a novel approach called VulEUT (Vulnerability Exploit Unit Test Generation), which combines vulnerability exploitation reachability analysis and LLM-based unit test generation. VulEUT is designed to automatically verify the exploitability of vulnerabilities in third-party libraries commonly used in client software projects. VulEUT first analyzes the client projects to determine the reachability of vulnerability conditions. And then, it leverages the Large Language Model (LLM) to generate unit tests for vulnerability confirmation. To evaluate the effectiveness of VulEUT, we collect 32 vulnerabilities from various third-party libraries and conduct experiments on 70 real client projects. Besides, we also compare our approach with two representative tools, i.e., TRANSFER and VESTA. Our results demonstrate the effectiveness of VulEUT, with 229 out of 292 generated unit tests successfully confirming vulnerability exploitation across 70 client projects, which outperforms baselines by 24%.
@@
The impact of vulnerabilities in TPLs has attracted increasing attention from both academia and industry. In recent years, automated tools have been introduced to detect and evaluate whether client projects are affected by vulnerabilities within the TPLs they depend on. For example, dependency-based approaches [6, 16, 17, 9] analyze dependency configuration files or gather compilation information through Software Composition Analysis techniques to identify vulnerable dependencies that may introduce security risks. Call graph-based approaches [18, 19, 20, 21] assess if projects invoke vulnerable code within TPLs, primarily focusing on the invocation of methods within the vulnerable code.
@@
We implement VulEUT, a tool capable of generating unit tests for client projects to automatically confirm whether TPL vulnerabilities are triggered.
@@ III-C3 Test Confirmation
This question primarily focuses on evaluating VulEUT’s effectiveness in vulnerability confirmation. We collect and analyze the number and accuracy of vulnerability confirmation tests generated by the VulEUT. We compare these results with those generated by TRANSFER and VESTA.
@@ IV-A Experimental Setup
Dataset. We collect 30 TPL vulnerabilities from publicly disclosed vulnerability databases such as CVE, covering various categories including XML data injection, remote code execution, and denial of service. For each collected vulnerability type, we collect client projects from GitHub as our experiment targets. We follow several criteria throughout the collection process: first, each vulnerability needs to have a corresponding publicly available Proof of Concept (PoC) code to ensure the confirmation of the vulnerability’s existence; second, client projects are written in Java, managed under the Maven [37] system, and are capable of being compiled and executed successfully; lastly, client projects depend on versions of TPL vulnerabilities, and based on this, we use call graph analysis to filter out method calls on vulnerable code. We exclude projects showing package dependencies on TPLs but do not utilize the vulnerable code. Finally, from the remaining project results, we select the first two client projects as experimental projects corresponding to each vulnerability type.
@@ IV-B RQ1: VulEUT Effectiveness
In the experiments, the VulEUT can complete code analysis and automatically generate prompts for unit tests within 30 seconds, showing its exceptional performance and potential for practical application. {mdframed}[backgroundcolor=gray!7] Answer to RQ1: VulEUT can generate unit tests for confirming vulnerability exploitation in client projects. It successfully generates 229 verifiable tests from 70 client projects, substantially outperforming the baseline tools (TRANSFER and VESTA), which demonstrates the effectiveness of our approach.
@@ IV-C RQ2: Ablation Study
[backgroundcolor=gray!7] Answer to RQ2: Method call path analysis, especially with parameter transfer analysis, significantly improved VulEUT performance. Prompt templates, particularly when combined with few-shot examples, considerably enhanced VulEUT performance, demonstrating the effectiveness of this approach.
@@ VI RELATED WORK
Security Issues of Third-party Libraries. Security issues related to third-party libraries have been a prominent focus of research [41, 15, 42, 43, 6, 44, 45, 46]. Kula et al. [2] conducted an investigation into the extent to which developers update dependencies in the Java ecosystem and how they respond to vulnerability notifications. The results showed that 81.5% of programs retained outdated dependencies without addressing vulnerabilities. Huang et al. [5] performed a risk analysis of 806 open-source projects and 544 security vulnerabilities within the Java ecosystem, focusing on usage risks and update risks. Zhan et al. [47] estimated that around 74.95% of the TPLs that contain vulnerabilities are widely utilized by other libraries. They provided practical recommendations for mitigating potential risks associated with maintaining third-party libraries, such as intelligent alerts and automated updates for outdated third-party libraries. With the booming of automated tools detecting and assessing the vulnerabilities in open-source projects, various kinds of problems are reported on the existing tools. For example, false alerts may be generated when using dependency analysis tools, and upgrading dependency versions may cause API-breaking due to bug fixing or code refactoring [9, 48, 49, 50]. Lacking existing tools [9, 49, 50] accessing the security threats of upstream vulnerabilities to downstream projects leads to a huge amount of projects being exposed to security threats. Mohayeji et al. [51] conducted an empirical study on the security of 978 JavaScript projects on GitHub with respect to third-party library dependency maintenance. They explored how developers responded to 4,195 security update recommendations made by a popular security robot called Dependabot. The results indicated that over half of the recommendations were accepted. Different from the aforementioned studies, our study focuses on evaluating whether vulnerabilities in these third-party libraries present real security risks to client projects, which aims to analyze a different problem.
@@
Vulnerability Reachability. Security issues related to TPLs have been a prominent focus of research [41, 15, 42, 43, 6, 44, 45, 46]. Ponta et al. [19] combined static and dynamic analysis tools to detect the reachability of vulnerable code, which overcomes their mutual limitations. Mir et al. [52] investigated the propagation patterns of vulnerabilities in the Maven ecosystem. They analyzed the distribution of vulnerabilities and the differences in vulnerability propagation at the package and method levels. Their approach relied on project dependencies, program call graphs, and conducted reachability analysis. The study emphasized the critical importance of considering vulnerability reachability at various granularity levels, particularly for security-focused applications. They suggested that graph-based analysis methods hold promise for future research. Wu et al. [9] conducted an empirical study on the impact of TPL vulnerabilities on software projects, analyzing CVEs, libraries, and project data. They examined vulnerability reachability and path complexity using call and control flow graphs but did not actively trigger these vulnerabilities in practice. Kang et al. proposed TRANSFER [14] to analyze the call relation of the methods in the client project and implied existing tools to generate a test case similar to the vulnerable witness test, which is considered a test case that could trigger the vulnerability. By executing the generated test case and manually checking the trigger condition of the vulnerability, TRANSFER accessed the reachability of the vulnerability precisely. Different from the aforementioned studies, our approach not only analyzes the reachability of vulnerability parameters between the client entry point and TPL vulnerability code but also generates unit test cases to confirm the presence of vulnerabilities. Our method checks the reachability of the vulnerabilities by generating a test case that covers the vulnerable method, adding an assertion to check the trigger of the vulnerability, and avoiding false alerts.
Fuzzing Techniques. Fuzzing techniques are also commonly employed to generate inputs for testing, aiming to uncover potential security vulnerabilities within projects. Kyriakos et al. [53] proposed an automated tool called FuzzGen, which utilizes whole-system analysis to infer the interface of a TPL and synthesizes a fuzzer specifically for that library, successfully uncovering multiple vulnerabilities in Debian and the Android Open Source Project. Domagoj et al. [54] introduced a new method named Directed Greybox Fuzzing (DGF), which efficiently generates inputs to reach specified target program locations, leading to the discovery of multiple security vulnerabilities and the assignment of CVE identifiers. Our approach differs from traditional fuzzing methods, which may struggle to generate meaningful inputs for exploiting vulnerabilities in certain scenarios such as JNDI injection. Instead, we extract domain-specific inputs from PoCs and incorporate them into test generation, enabling the generated tests to exploit vulnerabilities in client projects.
LLM-based Test Generation. There has been a growing body of research on the use of LLMs in test generation [31, 28, 29, 30, 55, 56]. Yuan et al. [30] explored ChatGPT’s capabilities in generating unit tests and introduced CHATTESTER, a tool that allows LLMs to understand source code for a focal method and generate a unit test for it. They used a validate-and-fix paradigm to validate the correctness of unit tests and automatically provided error information to the LLM to correct errors in unit tests, thereby improving the quality of unit tests generated by LLMs. Lemieux et al. [28] combined LLMs with traditional search-based test generation methods. They initially used a mutation algorithm to generate unit tests and, when they reached a coverage stall state, analyzed the currently generated tests and leveraged LLMs’ code generation capabilities to generate unit tests, resulting in higher test coverage. Wang et al. [57] fine-tuned the CodeT5 model to perform test completion for any targeted method under test and the test signature written by human developers. Deng et al. [58] complied a fuzzier with LLM to fuzz deep learning libraries by replacing the seed generation step and test cases mutation step in the traditional genetic algorithms with infilling large language model. which detected 65 bugs in Pytorch and TensorFlow and is the first work to demonstrate that LLMs can perform both generation-based and mutation-based fuzzing studied for decades. Unlike previous research focused solely on unit test generation tasks, our approach, VulnEUT, combines vulnerability analysis with LLM’s test generation capabilities to verify the exploitability of TPL vulnerabilities in client projects, enhancing the tool’s detection performance.

### v2

# Vulnerability-Triggering Test Case Generation from Third-Party Libraries

@@ Abstract
Open-source third-party libraries are widely used in software development. These libraries offer substantial advantages in terms of time and resource savings. However, a significant concern arises due to the publicly disclosed vulnerabilities within these libraries. Existing automated vulnerability detection tools often suffer from false positives and fail to accurately assess the propagation of inputs capable of triggering vulnerabilities from client projects to vulnerable code in libraries. In this paper, we propose a novel approach called VulEUT (Vulnerability Exploit Unit Test Generation), which combines vulnerability exploitation reachability analysis and LLM-based unit test generation. VulEUT is designed to automatically verify the exploitability of vulnerabilities in third-party libraries commonly used in Java client software projects. VulEUT first analyzes the client projects to determine the reachability of vulnerability conditions. And then, it leverages the Large Language Model (LLM) to generate unit tests for vulnerability confirmation. To evaluate the effectiveness of VulEUT, we collect 32 vulnerabilities from various third-party libraries and conduct experiments on 70 real client projects. Besides, we also compare our approach with two representative tools, i.e., TRANSFER and VESTA. Our results demonstrate the effectiveness of VulEUT, with 229 out of 292 generated unit tests successfully confirming vulnerability exploitation across 70 client projects, which outperforms baselines by 24%.
@@
The impact of vulnerabilities in TPLs has attracted increasing attention from both academia and industry. In recent years, automated tools have been introduced to detect and evaluate whether client projects are affected by vulnerabilities within the TPLs they depend on. For example, dependency-based approaches [6, 16, 17, 9] analyze dependency configuration files or gather compilation information through Software Composition Analysis techniques to identify vulnerable dependencies that may introduce security risks. Call graph-based approaches [18, 19, 20, 21, 19] assess if projects invoke vulnerable code within TPLs, primarily focusing on the invocation of methods within the vulnerable code.
@@
We implement VulEUT, a tool capable of generating unit tests for client projects to automatically confirm whether TPL vulnerabilities are triggered. Our tool is available on our website [32].
@@ IV-A Experimental Setup
Dataset. We collect 30 TPL vulnerabilities from publicly disclosed vulnerability databases such as CVE, covering various categories including XML data injection, remote code execution, and denial of service. For each collected vulnerability type, we collect client projects from GitHub as our experiment targets. We follow several criteria throughout the collection process: first, each vulnerability needs to have a corresponding publicly available Proof of Concept (PoC) code to ensure the confirmation of the vulnerability’s existence; second, client projects are written in Java, managed under the Maven [38] system, and are capable of being compiled and executed successfully; lastly, client projects depend on versions of TPL vulnerabilities, and based on this, we use call graph analysis to filter out method calls on vulnerable code. We exclude projects showing package dependencies on TPLs but do not utilize the vulnerable code. Finally, from the remaining project results, we select the first two client projects as experimental projects corresponding to each vulnerability type. These CVEs, TPLs, and client project lists, among other detailed data, are publicly available at [39].
@@ IV-B RQ1: VulEUT Effectiveness
In the experiments, the VulEUT can complete code analysis and automatically generate prompts for unit tests within 30 seconds, showing its exceptional performance and potential for practical application.
@@ VI RELATED WORK
Vulnerability Reachability. Security issues related to TPLs have been a prominent focus of research [43, 15, 44, 45, 6, 46, 47, 48]. Mir et al. [49] investigated the propagation patterns of vulnerabilities in the Maven ecosystem. They analyzed the distribution of vulnerabilities and the differences in vulnerability propagation at the package and method levels. Their approach relied on project dependencies, program call graphs, and conducted reachability analysis. The study emphasized the critical importance of considering vulnerability reachability at various granularity levels, particularly for security-focused applications. They suggested that graph-based analysis methods hold promise for future research. Kang et al. proposed TRANSFER [14] to analyze the call relation of the methods in the client project and implied existing tools to generate a test case similar to the vulnerable witness test, which is considered a test case that could trigger the vulnerability. By executing the generated test case and manually checking the trigger condition of the vulnerability, TRANSFER accessed the reachability of the vulnerability precisely. Different from the aforementioned studies, our approach not only analyzes the reachability of vulnerability parameters between the client entry point and TPL vulnerability code but also generates unit test cases to confirm the presence of vulnerabilities. Our method checks the reachability of the vulnerabilities by generating a test case that covers the vulnerable method, adding an assertion to check the trigger of the vulnerability, and avoiding false alerts.
@@ VII Conclusion and Future Work
Acknowledgment
This work was supported by the National Key R&D Program of China (No. 2024YFB4506400) and Ningbo Natural Science Foundation (No. 2023J292).